One son, one mother, one father, total age is 100yrs ,, son is half age of mother,, and mother is -5yrs of father.. Pls help me to calculate the age of son,,,,, the age of mother,,,,, the age of father
Answer by fractalier(6550)

Call their ages s, m, and f. Then we have
s + m + f = 100
s = (1/2)m
m = f - 5 (or f = m + 5)
Now substitute into the first equation and get
(1/2)m + m + m+5 = 100
(5/2)m + 5 = 100
(5/2)m = 95
m = 38
That means
s = 19 and
f = 43